# On the ring of cooperations for Real Hermitian K-theory

### v1

@@ 7.3. $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right)$
in Atiyah–Hirzebruch filtration 0 is entirely $h_{1}$-torsion, hence survives to $E_{3}$. The $d_{3}$-differential between Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 2\,(4)$ and Atiyah–Hirzebruch filtration 3 $cw\equiv 1\,(4)$ can be recovered from the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1))$ (see Fig. 17). Notice that there can be no differentials whose target are the $h_{0}$-towers $\bigoplus_{j=1}^{2k}\Sigma^{4j-4,4\lfloor j/2\rfloor-2}H$ for degree reasons. This determines the contributions to the $E_{\infty}$-page in coweight $cw\equiv 2\,(4)$ coming from Atiyah–Hirzebruch filtration 0.
The $d_{1}$-differential between Atiyah–Hirzebruch filtrations 2 and 3 in coweight $cw\equiv 1\,(4)$ can be recovered from the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1))$ (see Fig. 11). Similarly, the $d_{2}$-differential between Atiyah–Hirzebruch filtrations 0 and 2 in coweight $cw\equiv 1\,(4)$ can be recovered from the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1))$ (See Fig. 14). For degree reasons, we have that $E_{3}=E_{\infty}$ on Atiyah–Hirzebruch filtration 2, so this determines the contributions to the $E_{\infty}$-page in coweight $cw\equiv 2\,(4)$ from Atiyah–Hirzebruch filtration 2.
The $d_{1}$-differential between Atiyah–Hirzebruch filtrations 2 and 3 kills all classes from Atiyah–Hirzebruch filtration 3 in coweight $cw\equiv 0\,(4)$ which support $h_{0}$-multiplication. In particular, this kills all $h_{0}$-towers in the summand $\bigoplus_{j=1}^{2k}\Sigma^{4j-4,r\lfloor(j+1)/2\rfloor-4}H$ and in the big flag $\mathcal{F}_{4i,i,2i}$. The $d_{3}$-differential between Atiyah–Hirzebruch filtrations 0 and 3 is determined by the formula $d_{3}(\rho[3])=(\tau h_{1})[0]$ established in Lemma 7.1. In particular, using the notation of Definition 6.4, we have
$d_{3}(\rho\cdot x_{4k}[3])=(\Sigma^{2i,i}\tau h_{1})[0].$
Linearity gives the rest of the differentials. Note that this is the differential which ensures that Atiyah–Hirzebruch filtration 0 is trivial in coweight $cw\equiv 1\,(4)$ on the $E_{\infty}$-page.
Thus, in coweight $cw\equiv 2\,(4)$, the $E_{\infty}$-page is given by the remaining $h_{0}$-towers from the big staircase $\Sigma^{2i,i}\mathcal{S}$ from Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 2\,(4)$; the $h_{0}$-towers which do not come from the big flag
$\bigoplus_{j=1}^{2k}\Sigma^{4j-4,4\lfloor j/2\rfloor-2}H$
from Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 2\,(4)$; the remaining $\mathbb{F}_{2}[\tau^{4}]$’s from Atiyah–Hirzebruch filtration 2 in coweight $cw\equiv 2\,(4)$; and the remaining submodule of the big flag $\mathcal{F}_{4i,i,2i}$ from Atiyah–Hirzebruch filtration 3 in coweight $cw\equiv 0\,(4)$. Base change to $\mathbb{C}$ and linearity give the remaining hidden extensions, which we discuss now.
@@
By inspection, we see that the hidden extensions assemble all of the remaining classes on the $E_{\infty}$-page into the big flag
@@
Figure 25. $Z_{5}^{\mathbb{R}}$ in $cw\equiv 2\,(4)$.
Coweight $cw\equiv 3\,(4)$
@@
Finally, the coweight $cw\equiv 3\,(4)$ portion of the spectral sequence consists of classes coming from Atiyah–Hirzebruch filtration 2 in coweight $cw\equiv 2\,(4)$ and from Atiyah–Hirzebruch filtration 3 in coweight $cw\equiv 1\,(4)$. The $d_{1}$-differential between Atiyah–Hirzebruch filtrations 2 and 3 kills all classes from Atiyah–Hirzebruch filtration 2 in coweight $cw\equiv 2\,(4)$ which are $h_{0}$-divisible. In particular, only the bases of the $h_{0}$-towers $\bigoplus_{j=1}^{2k}\Sigma^{4j-4,4\lfloor j/2\rfloor-2}H$ survive, to the $E_{2}$-page:
$\bigoplus_{j=1}^{2k}\Sigma^{4j-4,4\lfloor j/2\rfloor-2}\mathbb{F}_{2}[\tau^{4}],$
@@
and only bases of the $h_{0}$-towers in the big staircase $\Sigma^{2i,i}\mathcal{S}$, as well as the classes that are not $h_{0}$-divisible, survive. The $d_{2}$-differential between Atiyah–Hirzebruch filtrations 0 and 2 kills all classes from Atiyah–Hirzebruch filtration 2 in coweight $cw\equiv 2\,(4)$ which support $h_{1}$-multiplication. The classes remaining from the $h_{0}$-towers that are not from the big staircase do not support $h_{1}$-multiplication, and neither do the two classes $b^{n}a$ and $b^{n}\rho a$ remaining from the staircases, so they survive to the $E_{3}$-page. For degree reasons, we have that $E_{3}=E_{\infty}$ on Atiyah–Hirzebruch filtration 2, so this determines the contributions to the $E_{\infty}$-page in coweight $cw\equiv 3\,(4)$ from Atiyah–Hirzebruch filtration 2.
@@
The $d_{1}$-differential between Atiyah–Hirzebruch filtrations 2 and 3 kills all classes from Atiyah–Hirzebruch filtration 3 in coweight $cw\equiv 1\,(4)$ which support $h_{0}$-multiplication. This can be recovered from the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1))$. The $d_{3}$-differential between Atiyah–Hirzebruch filtrations 3 and 0 is only nonzero on the bottom left corner of the remnants of the diamond. This determines the contributions to the $E_{\infty}$-page in coweight $cw\equiv 3\,(4)$ from Atiyah–Hirzebruch filtration 3.
@@
Thus, in coweight $cw\equiv 1\,(4)$, the $E_{\infty}$-page is given by the segments $T$ coming from Atiyah–Hirzebruch filtrations 2 and 3 and the bases of the $h_{0}$-towers from Atiyah–Hirzebruch filtration 2. These segments are connected by hidden extensions in the following way. Let $\Sigma^{2i,i}(h_{1}(\tau h_{1}))[3]$ and $\Sigma^{2i,i}\rho h_{1}(\tau h_{1})[3]$ be the classes remaining from Atiyah–Hirzebruch filtration 3 in lowest Adams filtration, and let $\Sigma^{2i,i}a[2]$ and $\Sigma^{2i,i}\rho a[2]$ be the classes remaining from Atiyah–Hirzebruch filtration 2 in lowest Adams filtration. Base change to $\mathbb{C}$ gives a hidden extension
@@
Linearity over $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ gives us that
@@
Multiplication by $b$ gives the same extensions throughout the $E_{\infty}$-page. Altogether, this gives the summand
@@
For degree reasons, the bases of the $h_{0}$-towers from Atiyah–Hirzebruch filtration 2 are $v_{1}$-torsion.
Figure 26. $Z_{5}^{\mathbb{R}}$ in $cw\equiv 3\,(4)$.
[…]
@@
It is possible that one can recover 7.10 by using the $\rho$-Bockstein spectral sequence and [CQ21, Lemma 3.36]. This spectral sequence has signature
$E_{1}=\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}\left(B_{0}^{\mathbb{C}}(1)^{\otimes i}\right)[\rho]\implies\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right).$
@@ 8. The ring of cooperations
In this section, we continue the inductive procedure and compute the groups $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(k))$. Then, we assemble our results to compute the ring of cooperations $\pi_{**}^{\mathbb{R}}(\text{kq}\otimes\text{kq})$.
8.1. $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(k))$
The goal of this section is to compute $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(k))$. We begin with the following lemma.
@@ 8.1. $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(k))$
There is an isomorphism of $\mathcal{A}(1)^{\vee}_{\mathbb{R}}$-comodules and $\textup{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-modules:
$\textup{Ext}_{\mathcal{A}(0)^{\vee}_{\mathbb{R}}}(B_{1}^{\mathbb{R}}(k))\cong\bigoplus_{j=0}^{k}\Sigma^{4j,2j}\mathbb{F}_{2}[\rho,\tau^{2},h_{0}]/(\rho h_{0}).$
@@
Let $C_{\mathcal{A}(0)^{\vee}}(B_{1}^{\mathbb{R}}(k))$ denote the cobar complex computing $\text{Ext}_{\mathcal{A}(0)^{\vee}_{\mathbb{R}}}(B_{1}^{\mathbb{R}}(k))$. This is a cosimplicial ring which takes the form
[…]
@@
$E_{1}=\text{Ext}_{\mathcal{A}(0)^{\vee}_{\mathbb{C}}}(B_{1}^{\mathbb{C}}(k))[\rho]\implies\text{Ext}_{\mathcal{A}(0)^{\vee}_{\mathbb{R}}}(B_{1}^{\mathbb{R}}(k)).$
The groups $\text{Ext}_{\mathcal{A}(0)^{\vee}_{\mathbb{C}}}(B_{1}^{\mathbb{C}}(k))$ were calculated by Culver-Quigley [CQ21, Remark 3.25]. Modulo $v_{1}$-torsion, we have
$\text{Ext}_{\mathcal{A}(0)^{\vee}_{\mathbb{C}}}(B_{1}^{\mathbb{C}}(k))\cong\bigoplus_{j=0}^{k}\Sigma^{4j,2j}\mathbb{M}^{\mathbb{C}}_{2}[h_{0}].$
In the case of $k=0$, we have that $B_{1}^{\mathbb{C}}(0)\cong\mathbb{M}_{2}^{\mathbb{C}}$, and this $\rho$-Bockstein spectral sequence was completely determined by Hill [Hil11, Theorem 3.1]. We can rewrite the $E_{1}$-page as
$E_{1}=\text{Ext}_{\mathcal{A}(0)^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})[\rho]=\mathbb{M}_{2}^{\mathbb{C}}[h_{0}][\rho],$
@@
$\text{Ext}_{\mathcal{A}(0)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})\cong\mathbb{F}_{2}[\rho,\tau^{2},h_{0}]/(\rho h_{0}).$
More generally, the $E_{1}$-page of the $\rho$-Bockstein takes the form
$E_{1}=\bigoplus_{j=0}^{k}\Sigma^{4j,2j}\mathbb{M}_{2}^{\mathbb{C}}[h_{0}][\rho],$
@@
for each $0\leq j\leq k$ on the classes $\tau$ in each $H$-tower in the decomposition of $\text{Ext}_{\mathcal{A}(0)^{\vee}_{\mathbb{C}}}(B_{1}^{\mathbb{C}}(k))$. The Liebniz rule immediately determines the end result:
$\text{Ext}_{\mathcal{A}(0)^{\vee}_{\mathbb{R}}}(B_{1}^{\mathbb{R}}(k))\cong\bigoplus_{j=0}^{k}\Sigma^{4j,2j}\mathbb{F}_{2}[\rho,\tau^{2},h_{0}]/(\rho h_{0}),$
@@
Figure 27. The group $\text{Ext}_{\mathcal{A}(0)^{\vee}_{\mathbb{R}}}(B_{1}^{\mathbb{R}}(k)).$ A black $\blacksquare$ denotes $\mathbb{F}_{2}[\rho,\tau^{2}]$, and a black $\bullet$ denotes $\mathbb{F}_{2}[\tau^{2}]$. A vertical black line represents multiplication by $h_{0}$.
We will now use the results of 7.10 and the short exact sequences of motivic Brown–Gitler comodules to culminate our algebraic Atiyah–Hirzebruch program. The following result should be compared with [Mah81, Prop 2.6], [BOSS19, Prop 3.3], and 5.6.
@@
There is an isomorphism of $\mathcal{A}(1)_{\mathbb{R}}^{\vee}$-comodules and $\textup{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-modules:
$\frac{\textup{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(k))}{v_{1}\textup{-torsion}}\cong\Sigma^{4k-4,2k-2}Z_{\mathbb{\alpha}(k)}^{\mathbb{R}}\oplus\bigoplus_{j=0}^{4k-8}\left(\Sigma^{4j,2j}H\oplus\Sigma^{4j,2j-2}H\right),$
where $\alpha(k)$ is the number of 1’s in the dyadic expansion of $k$.

### v2

@@ 7.3. $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right)$
which do not come from the big staircase from Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 2\,(4)$; the remaining $\mathbb{F}_{2}[\tau^{4}]$’s from Atiyah–Hirzebruch filtration 2 in coweight $cw\equiv 1\,(4)$; and the remaining submodule of the big flag $\mathcal{F}_{4i,i,2i}$ from Atiyah–Hirzebruch filtration 3 in coweight $cw\equiv 0\,(4)$. Base change to $\mathbb{C}$ and linearity give the remaining hidden extensions, which we discuss now.
@@
By inspection, we see that the hidden extensions assemble all of the remaining classes on the $\textup{E}_{\infty}$-page into the big flag
@@
The reader may benefit from consulting the chart in Figure 7.10.
As an example, Figure 7.17 depicts the coweight $cw\equiv 2\,(4)$ portion of $Z_{5}^{\mathbb{R}}$. In this case, we have:
  • the $h_{0}$-towers in stems 0 and 4 come from the $h_{0}$-towers in Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 2\,(4)$ which do not come from the big staircase;
  • the big flag $\mathcal{F}_{16,4,8}$ consists of the $h_{0}$-towers from the big staircase from Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 2\,(4)$, the remaining $\mathbb{F}_{2}[\tau^{4}]$’s from Atiyah–Hirzebruch filtration 2 in coweight $cw\equiv 1\,(4)$, and the remaining submodule of the big flag from Atiyah–Hirzebruch filtration 3 in coweight $cw\equiv 0\,(4)$, together with hidden extensions.
Figure 7.17. $Z_{5}^{\mathbb{R}}$ in $cw\equiv 2\,(4)$.
$\textup{E}_{\infty}$-coweight $cw\equiv 3\,(4)$
@@
Finally, the coweight $cw\equiv 3\,(4)$ portion of the spectral sequence consists of classes coming from Atiyah–Hirzebruch filtration 2 in coweight $cw\equiv 2\,(4)$ and from Atiyah–Hirzebruch filtration 3 in coweight $cw\equiv 1\,(4)$. The $d_{1}$-differential between Atiyah–Hirzebruch filtrations 2 and 3 kills all classes from Atiyah–Hirzebruch filtration 2 in coweight $cw\equiv 2\,(4)$ which are $h_{0}$-divisible. In particular, only the bases of the $h_{0}$-towers $\bigoplus_{j=1}^{2k}\Sigma^{4j-4,4\lfloor j/2\rfloor-2}H$ survive, to the $\textup{E}_{2}$-page:
(7.8) $\bigoplus_{j=1}^{2k}\Sigma^{4j-4,4\lfloor j/2\rfloor-2}\mathbb{F}_{2}[\tau^{4}],$
@@
and only bases of the $h_{0}$-towers in the big staircase $\Sigma^{2i,i}\mathcal{S}$, as well as the classes that are not $h_{0}$-divisible, survive (see Figure 7.3 for the differential on the big staircase summands and Figure 7.15 for the differential on the $h_{0}$-towers). The $d_{2}$-differential between Atiyah–Hirzebruch filtrations 0 and 2 kills all classes from Atiyah–Hirzebruch filtration 2 in coweight $cw\equiv 2\,(4)$ which support $h_{1}$-multiplication. The classes in (7.8) do not support $h_{1}$-multiplication, and neither do the classes $b^{n}a$ and $b^{n}\rho a$ remaining from the staircases (see Figure 7.6), so they survive to the $\textup{E}_{3}$-page. For degree reasons, we have that $\textup{E}_{3}=\textup{E}_{\infty}$ on Atiyah–Hirzebruch filtration 2, so this determines the contributions to the $\textup{E}_{\infty}$-page in coweight $cw\equiv 3\,(4)$ from Atiyah–Hirzebruch filtration 2.
@@
The $d_{1}$-differential between Atiyah–Hirzebruch filtrations 2 and 3 kills all classes from Atiyah–Hirzebruch filtration 3 in coweight $cw\equiv 1\,(4)$ which support $h_{0}$-multiplication. This can be recovered from the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1))$ (see Figure 7.2). The $d_{3}$-differential between Atiyah–Hirzebruch filtrations 3 and 0 is only nonzero on the bottom left corner of the remnants of the diamond (see Figure 7.8). This determines the contributions to the $\textup{E}_{\infty}$-page in coweight $cw\equiv 3\,(4)$ from Atiyah–Hirzebruch filtration 3.
@@
Thus, in coweight $cw\equiv 1\,(4)$, the $\textup{E}_{\infty}$-page is given by the segments $T$ coming from Atiyah–Hirzebruch filtrations 2 and 3 and the bases of the $h_{0}$-towers from Atiyah–Hirzebruch filtration 2. These segments are connected by hidden extensions in the following way. Let $\Sigma^{2i,i}(h_{1}(\tau h_{1}))[3]$ and $\Sigma^{2i,i}\rho h_{1}(\tau h_{1})[3]$ be the classes remaining from Atiyah–Hirzebruch filtration 3 in lowest Adams filtration, and let $\Sigma^{2i,i}a[2]$ and $\Sigma^{2i,i}\rho a[2]$ be the classes remaining from Atiyah–Hirzebruch filtration 2 in lowest Adams filtration. Base change to $\mathbb{C}$ gives a hidden extension
@@
Linearity over $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ gives us that
@@
Multiplication by $b$ gives the same extensions throughout the $\textup{E}_{\infty}$-page. Altogether, this gives the summand
@@
For degree reasons, the bases of the $h_{0}$-towers from Atiyah–Hirzebruch filtration 2 are $v_{1}$-torsion. The reader may benefit from consulting the chart in Figure 7.11.
As an example, Figure 7.18 depicts the coweight $cw\equiv 3\,(4)$ portion of $Z_{5}^{\mathbb{R}}$. In this case, we have only a $\mathcal{D}\langle 1\rangle$ in stem 13 which consists of the segments from Atiyah–Hirzebruch filtration 2 in coweight $cw\equiv 2\,(4)$ and from Atiyah–Hirzebruch filtration 3 in coweight $cw\equiv 1\,(4)$, together with hidden extensions.
Figure 7.18. $Z_{5}^{\mathbb{R}}$ in $cw\equiv 3\,(4)$.
Thus, we have shown that for $i=4k+1$, $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right)/v_{1}\text{-torsion}$ is given by the direct sum of columns of Table 5. The other congruence classes are proved in a similar fashion. ∎
@@
It is possible that one can recover 7.6 by using the $\rho$-Bockstein spectral sequence and [17, Lemma 3.36]. This spectral sequence has signature
$\textup{E}_{1}=\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}\left(B_{0}^{\mathbb{C}}(1)^{\otimes i}\right)[\rho]\implies\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right).$
@@ 8. The ring of cooperations
In this section, we continue the inductive procedure and compute the groups $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(k))$. Then, we assemble our results to compute the ring of cooperations $\pi_{**}^{\mathbb{R}}(\text{kq}\otimes\text{kq})$. As an application, we determine the structure of the $\text{E}_{1}$-page of the $\mathbb{R}$-motivic kq-resolution.
8.1. $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(k))$
The goal of this section is to compute $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(k))$. We begin with the following lemma.
@@ 8.1. $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(k))$
There is an isomorphism of $\mathcal{A}(1)^{\vee}_{\mathbb{R}}$-comodules and $\textup{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-modules:
$\textup{Ext}^{***}_{\mathcal{A}(0)^{\vee}_{\mathbb{R}}}(B_{1}^{\mathbb{R}}(k))\cong\bigoplus_{j=0}^{k}\Sigma^{4j,2j}\mathbb{F}_{2}[\rho,\tau^{2},h_{0}]/(\rho h_{0}).$
@@
Let $C_{\mathcal{A}(0)_{\mathbb{R}}^{\vee}}(B_{1}^{\mathbb{R}}(k))$ denote the cobar complex computing $\text{Ext}^{***}_{\mathcal{A}(0)^{\vee}_{\mathbb{R}}}(B_{1}^{\mathbb{R}}(k))$. This is a cosimplicial ring which takes the form
$B_{1}^{\mathbb{R}}(k)$$\mathcal{A}(0)^{\vee}_{\mathbb{R}}\otimes B_{1}^{\mathbb{R}}(k)$$\mathcal{A}(0)^{\vee}_{\mathbb{R}}\otimes\mathcal{A}(0)^{\vee}_{\mathbb{R}}\otimes B_{1}^{\mathbb{R}}(k)$$\cdots$}
@@
$\textup{E}_{1}=\text{Ext}^{***}_{\mathcal{A}(0)^{\vee}_{\mathbb{C}}}(B_{1}^{\mathbb{C}}(k))[\rho]\implies\text{Ext}^{***}_{\mathcal{A}(0)^{\vee}_{\mathbb{R}}}(B_{1}^{\mathbb{R}}(k)).$
The groups $\text{Ext}^{***}_{\mathcal{A}(0)^{\vee}_{\mathbb{C}}}(B_{1}^{\mathbb{C}}(k))$ were calculated by Culver-Quigley [17, Remark 3.25]. Modulo $v_{1}$-torsion, we have
$\text{Ext}^{***}_{\mathcal{A}(0)^{\vee}_{\mathbb{C}}}(B_{1}^{\mathbb{C}}(k))\cong\bigoplus_{j=0}^{k}\Sigma^{4j,2j}\mathbb{M}^{\mathbb{C}}_{2}[h_{0}].$
In the case of $k=0$, we have that $B_{1}^{\mathbb{C}}(0)\cong\mathbb{M}_{2}^{\mathbb{C}}$, and this $\rho$-Bockstein spectral sequence was completely determined by Hill [26, Theorem 3.1]. We can rewrite the $\textup{E}_{1}$-page as
$\textup{E}_{1}=\text{Ext}^{***}_{\mathcal{A}(0)^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})[\rho]=\mathbb{M}_{2}^{\mathbb{C}}[h_{0}][\rho],$
@@
$\text{Ext}^{***}_{\mathcal{A}(0)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})\cong\mathbb{F}_{2}[\rho,\tau^{2},h_{0}]/(\rho h_{0}).$
More generally, the $\textup{E}_{1}$-page of the $\rho$-Bockstein takes the form
$\textup{E}_{1}=\bigoplus_{j=0}^{k}\Sigma^{4j,2j}\mathbb{M}_{2}^{\mathbb{C}}[h_{0}][\rho],$
@@
for each $0\leq j\leq k$ on the classes $\tau$ in each $H$-tower in the decomposition of $\text{Ext}^{***}_{\mathcal{A}(0)^{\vee}_{\mathbb{C}}}(B_{1}^{\mathbb{C}}(k))$. The Liebniz rule immediately determines the end result:
$\text{Ext}^{***}_{\mathcal{A}(0)^{\vee}_{\mathbb{R}}}(B_{1}^{\mathbb{R}}(k))\cong\bigoplus_{j=0}^{k}\Sigma^{4j,2j}\mathbb{F}_{2}[\rho,\tau^{2},h_{0}]/(\rho h_{0}),$
@@
Figure 8.1. The group $\text{Ext}^{***}_{\mathcal{A}(0)^{\vee}_{\mathbb{R}}}(B_{1}^{\mathbb{R}}(k)).$ A black $\blacktriangle$ denotes $\mathbb{F}_{2}[\rho,\tau^{2}]$, and a black $\bullet$ denotes $\mathbb{F}_{2}[\tau^{2}]$. A vertical black line represents multiplication by $h_{0}$.
We will now combine the results of 7.6 and the short exact sequences of motivic Brown–Gitler comodules from Proposition 2.11. The following result should be compared with [41, Prop 2.6], [10, Prop 3.3], and 5.5.
@@
There is an isomorphism of $\mathcal{A}(1)_{\mathbb{R}}^{\vee}$-comodules and $\textup{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-modules:
$\frac{\textup{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(k))}{v_{1}\textup{-torsion}}\cong\Sigma^{4k-4,2k-2}Z_{\mathbb{\alpha}(k)}^{\mathbb{R}}\oplus\bigoplus_{j=0}^{4k-8}\left(\Sigma^{4j,2j}H\oplus\Sigma^{4j,2j-2}H\right),$
where $\alpha(k)$ is the number of 1’s in the dyadic expansion of $k$. There are $\tau^{2}$-extensions between $\Sigma^{4j,2j}H$ and $\Sigma^{4j,2j-2}H.$